In a two-tailed hypothesis test situation, the test statistic is determined to be t = -2.032. The sample size is 35. The p-value for this test is
Answer #1

Since , the test is two-tailed test.

d.f.=degrees of freedom=n-1=35-1=34

The p-value is ,

The Excel function for p-value is , =TDIST(2.032,34,2)